# Differentially private federated learning for localized control of infectious disease dynamics

### v1

@@ Abstract
In times of epidemics, swift reaction is necessary to mitigate epidemic spreading. For this reaction, localized approaches have several advantages, limiting necessary resources and reducing the impact of interventions on a larger scale. However, training a separate machine learning (ML) model on a local scale is often not feasible due to limited available data. Centralizing the data is also challenging because of its high sensitivity and privacy constraints. In this study, we consider a localized strategy based on the German counties and communities managed by the related local health authorities (LHA). For the preservation of privacy to not oppose the availability of detailed situational data, we propose a privacy-preserving forecasting method that can assist public health experts and decision makers. ML methods with federated learning (FL) train a shared model without centralizing raw data. Considering the counties, communities or LHAs as clients and finding a balance between utility and privacy, we study a FL framework with client-level differential privacy (DP). We train a shared multilayer perceptron on sliding windows of recent case counts to forecast the number of cases in the future, while clients exchange only norm-clipped updates and the server aggregated updates with DP noise. We evaluate the approach on COVID-19 data on county-level during two phases—November 2020 and March 2022 (Omicron). As expected, very strict privacy ($\varepsilon\leq 0.5$) yields unstable, unusable forecasts. At a moderately strong but still privacy-preserving level ($\varepsilon=2$), the DP model closely approaches the non-DP model: $R^{2}\approx 0.94$ (vs. 0.95) and mean absolute percentage error (MAPE) $\approx 26\%$ in November 2020; $R^{2}\approx 0.88$ (vs. $0.93$) and MAPE $\approx 21\%$ in March 2022. Overall, client-level DP-FL can deliver useful county-level predictions with strong privacy guarantees, and viable privacy budgets depend on epidemic phase, allowing privacy-compliant collaboration among health authorities for local forecasting.
@@ Introduction
In epidemic emergencies or situations of seasonally peaking endemic pathogens, public health experts and decision makers can benefit from predictions of infectious disease dynamics to proactively react to upcoming challenges. For prediction tasks, mechanistic mathematical models as well as ML models have proven to provide reliable outcomes. While for mechanistic metapopulation [1, 2], agent-based [3, 4, 5] or combined differential equation- and ML-based [6, 7] approaches, new situations might require the adaptations of existing mechanisms, many ML models are developed in a purely data-driven fashion, reducing the number of manual interventions cycles.
The traditional approach to training machine learning models typically requires centralizing data from multiple parties. This creates serious privacy risks when handling sensitive information such as medical records, which often cannot be shared directly. In infectious disease contexts, these records include precise dates of symptom onset and assumed transmission events, maintained by physicians, hospitals, or LHAs, and are critical for predicting future cases. While aggregation and anonymization can reduce exposure of sensitive information, they do not eliminate the risk of re-identification and often result in substantial information loss. Ensuring privacy-preserving approaches is therefore essential when working with such data.
@@
However, sharing gradients during collaborative learning can still reveal sensitive information from the training data of individual parties. Advanced attacks have demonstrated that adversaries can infer specific data records or group characteristics from intercepted model updates [11, 12, 13, 14, 15, 16, 17]. In extreme cases, they can reconstruct entire training samples from the gradients [16, 17]. Therefore, merely avoiding the sharing of actual training data can not completely avoid privacy leakage.
@@
Initial works on calibrating noise to sensitivity [18] led to an extensive work of algorithmic foundations of Differential Privacy [19]. DP has emerged as a preeminent framework for enhancing data privacy, countering various privacy attacks associated with FL. DP provides strong privacy assurances by ensuring that the behavior of the aggregated model does not depend significantly on any single client’s data. Instead, it captures commonalities across all contributions. Implementing DP involves adding Gaussian noise to the aggregated update, thereby masking participant data specifics while preserving the overall learning utility. In 2021, the authors of [20] had observed a substantial increase in the use of DP in health research publications and, among others, the authors of [21, 22, 23, 24] developed a framework for privacy-preserving FL. In the context of the COVID-19 pandemic, the authors of [25] raised the awareness for a potential distortion of COVID-19 death rates when using DP with too small dataset groupings. During the COVID-19 pandemic, more than 1.9 million COVID-19 cases across Europe have been integrated in the DataSHIELD (Data Aggregation Through Anonymous Summary-statistics from Harmonised Individual-levEL Databases) platform [26], to allow for FL approaches to derive potential strategies against COVID-19 [27]. Among other things, DataSHIELD has already been used for other infectious diseases, e.g., to study malaria intervention in Mozambique [28].
In order to report dynamics and allow predictions by simultaneously addressing privacy concerns, COVID-19 case counts had been aggregated and reported on county-level throughout the pandemic in Germany [29]. We use these official aggregates as a proof-of-concept dataset to evaluate our method, which is intended for settings where finer-grained, potentially sensitive data remain with the LHAs. While aggregation reduces disclosure risk, it does not provide formal privacy guarantees [30]. In this study, we propose a differentially private federated learning approach designed to improve the prediction of new cases of a circulating pathogen, while providing theoretical guarantees for the protection of each participants information (client-level DP). While we validate our approach on published data on county-level, we suggest to explore the potential of localized predictions on a finer spatial scale when available.
@@
Our objective is time‐series forecasting of local case counts using data which is reported on local level. We construct supervised learning examples by using a historical window of size $H=10$ days over each administrative unit’s time series and predicting the case count $P=7$ days ahead. Let $c_{k,t}$ denote the raw, unnormalized number of daily reported cases for geographic unit $k$, $k=1,2,3,\ldots$ on day $t$. Therefore, the input vector
@@
is the case count $P$ days into the future. Missing reports are handled as zero reported cases. From our dataset, we generate multiple input-label pairs for each administrative unit. Each input-label pair consists of an input window of $H$ consecutive days and a target value $P$ days ahead.
For model evaluation, we randomly select 10% of the generated input-label pairs as test set and the remaining 90% as training set. After splitting, we remove the location identifier from the feature vectors since they should not influence the model in any way.
@@
We use PyTorch [31] to build a fully connected multilayer perceptron (MLP) as our prediction model. The architecture is consistent across both spatial resolutions, community- and county-based; see the Results section for details on the datasets used. The input to the network is a vector of length $H$ containing the $H$ most recent daily case counts; with zero reported cases counting as daily case count. This vector is propagated through three fully connected hidden layers of sizes 128, 64, and 32 neurons, each followed by a rectified linear unit (ReLU) activation function. A final linear output unit produce the predicted case count $P$ days ahead. We use the mean squared error (MSE) as the loss function, which is minimized by the Adam optimizer with a learning rate of 0.001.
@@
The core methodology employs FL with client-level DP, summarized in Algorithm 1. All hyperparameters are additionally explained in Table 1. The training process basically follows four steps:
@@
1Server:
2 Initialize common model ${\mathbf{w}}_{0}$
3 for $i=1$ to $T_{\mathsf{cl}}$ do
4   Select a subset $\mathbb{K}$ of clients randomly with probability $q$
5   Initialize list of updates $U=\mathbf{[}\,\mathbf{]}$
6   for each client $k\in\mathbb{K}$ do
7     $\Delta\widehat{{\mathbf{w}}}_{i}^{k}=\mathbf{Client}_{k}({\mathbf{w}}_{i-1})$
8     Insert $\Delta\widehat{{\mathbf{w}}}_{i}^{k}$ into $U$
9
10     end for
11    $\Delta\overline{{\mathbf{w}}}_{i}=\frac{1}{m}\sum_{\Delta\widehat{{\mathbf{w}}}\in U}\Delta\widehat{{\mathbf{w}}}$
12     if DP enabled then
13       Calculate noise multiplier $c$ using an RDP accountant for $(\varepsilon,\delta,q,T_{\mathsf{cl}},E)$
14       $\sigma=(S\cdot c)/m$
15       Draw $\xi_{i}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$
16       ${\mathbf{w}}_{i}={\mathbf{w}}_{i-1}+\Delta\overline{{\mathbf{w}}}_{i}+\xi_{i}$
17
18       else
19         ${\mathbf{w}}_{i}={\mathbf{w}}_{i-1}+\Delta\overline{{\mathbf{w}}}_{i}$
20
21         end if
22
23         end for
24        $\mathbf{Client}_{k}({\mathbf{w_{i-1}}})$:
25         $\widetilde{{\mathbf{w}}}_{i}={\mathbf{w}}_{i-1}$
26         Train $\widetilde{{\mathbf{w}}}_{i}$ on local data $D_{k}$ for $E$ epochs using Adam ($\eta=0.001$)
27         $\Delta{\mathbf{w}}_{i}^{k}=\widetilde{{\mathbf{w}}}_{i}-{\mathbf{w}}_{i-1}$
@@
29         Output: $\Delta\widehat{{\mathbf{w}}}_{i}^{k}$
@@ Results
We use official reported county-level COVID-19 case data from the LHAs transferred to the Robert Koch Institute (RKI) [29] and accessed via the memilio-epidata Python package [33]. We consider two different pandemic stages, first, late-stage as of March 2022 and, second, early stage as of November 2020. For each county and day, we obtain the newly confirmed infections plus estimated recoveries and deaths reported with the (assumed) day of infection. This data covers all 400 German counties. While the published data is stratified on county-level, the counties can, again, be divided by communities, summing up to a total of 10,786 communities in Germany. In Fig 1, Panel A, shows the population distribution across the counties, highlighting the variation in population sizes. The median county population is 147,524, while the mean population is 200,525, as several counties have a substantial larger population. The largest county by population has 3,685,265 inhabitants (Berlin, after merging all Berlin districts), and the total population across all counties is approximately 83.6 million. This heterogeneity in county sizes presents a particular challenge for our FL approach. Counties with larger populations typically report higher absolute case numbers. In addition, we show in Panel B the distribution of population sizes on community-level.
@@
Fig 2: Spatial resolution of German counties and communities and community-based case data. Panel A presents two maps shown side by side. The left map presents North Rhine-Westphalia (NRW) stratified into counties (thick black boundaries) and communities (thin gray boundaries), highlighting county ”Coesfeld” in red. The right map displays the stratification of Germany into federal states (thick black boundaries) and counties (thin gray boundaries), highlighting the federal state NRW in orange. Panel B shows the obtained community-based dataset for ”Coesfeld” for March 2022, with the county-level aggregate shown in red and individual community trajectories in various colors. The left map from Panel A is using geodata “Verwaltungsgebiete 1:250 000 Stand $01.$ (VG250 01.01.)” from https://gdz.bkg.bund.de, copyright; GeoBasis-DE / BKG 2021, license dl-de/by-2–0, see https://www.govdata.de/dl-de/by-2-0.

### v2

@@ Abstract
In times of epidemics, swift reaction is necessary to mitigate epidemic spreading. For this reaction, localized approaches have several advantages, limiting necessary resources and reducing the impact of interventions on a larger scale. However, training a separate machine learning (ML) model on a local scale is often not feasible due to limited available data. Centralizing the data is also challenging because of its high sensitivity and privacy constraints. In this study, we consider a localized strategy based on the German counties and communities managed by the related local health authorities (LHA). For the preservation of privacy to not oppose the availability of detailed situational data, we propose a privacy-preserving forecasting method that can assist public health experts and decision makers. ML methods with federated learning (FL) train a shared model without centralizing raw data. Considering the counties, communities or LHAs as clients and finding a balance between utility and privacy, we study a FL framework with client-level differential privacy (DP). We train a shared multilayer perceptron on sliding windows of recent case counts to forecast the number of cases in the future, while clients exchange only norm-clipped updates and the server aggregated updates with DP noise. We evaluate the approach on COVID-19 data on county-level during two phases-November 2020 and March 2022 (Omicron). As expected, very strict privacy ($\varepsilon\leq 0.5$) yields unstable, unusable forecasts. At a moderately strong but still privacy-preserving level ($\varepsilon=2$), the DP model closely approaches the non-DP model: $R^{2}\approx 0.94$ (vs. 0.95) and mean absolute percentage error (MAPE) $\approx 26\%$ in November 2020; $R^{2}\approx 0.88$ (vs. $0.93$) and MAPE $\approx 21\%$ in March 2022. Overall, client-level DP-FL can deliver useful county-level predictions with strong privacy guarantees, and viable privacy budgets depend on epidemic phase, allowing privacy-compliant collaboration among health authorities for local forecasting.
@@ Introduction
In epidemic emergencies or situations of seasonally peaking endemic pathogens, public health experts and decision makers can benefit from predictions of infectious disease dynamics to proactively react to upcoming challenges. For prediction tasks, mechanistic mathematical models as well as ML models have proven to provide reliable outcomes. While for mechanistic population [1, 2, 3], metapopulation [4, 5], agent-based [6, 7, 8] or combined differential equation- and ML-based [9, 10] approaches, new situations might require the adaptations of existing mechanisms, many ML models are developed in a purely data-driven fashion, reducing the number of manual interventions cycles. In particular, classical simulation or forecasting studies [7, 1, 2, 3, 8] mostly used publicly available datasets and did not consider the aspect of privacy-preservation.
The traditional approach to training machine learning models typically requires centralizing data from multiple parties. This creates serious privacy risks when handling sensitive information such as medical records, which often cannot be shared directly. In infectious disease contexts, these records include precise dates of symptom onset (essential for estimating reporting delays), assumed transmission events, and detailed demographics maintained by hospitals or LHAs. These features are critical for predicting future cases but are removed from public datasets to preserve anonymity. While aggregation reduces disclosure risk, it removes these valuable signals and does not eliminate the risk of re-identification [11]. Ensuring privacy-preserving approaches is therefore essential when working with such data.
@@
However, sharing gradients during collaborative learning can still reveal sensitive information from the training data of individual parties. Advanced attacks have demonstrated that adversaries can infer specific data records or group characteristics from intercepted model updates [15, 16, 17, 18, 19, 20, 21, 22]. In extreme cases, they can reconstruct entire training samples from the gradients [20, 21]. Therefore, merely avoiding the sharing of actual training data can not completely avoid privacy leakage.
@@
Initial works on calibrating noise to sensitivity [23] led to an extensive work of algorithmic foundations of Differential Privacy [24]. DP has emerged as a preeminent framework for enhancing data privacy, countering various privacy attacks associated with FL. DP provides strong privacy assurances by ensuring that the behavior of the aggregated model does not depend significantly on any single client’s data. Instead, it captures commonalities across all contributions. In the context of federated learning, several foundational works have combined $(\varepsilon,\delta)$-DP with standard federated learning algorithms that aggregate client updates on a central server [25, 26, 27], or with local differential privacy mechanisms in which each client perturbs its model update before sending it to the server [28]. An overview of the federated learning landscape, including privacy preservation, is given by [29]. Implementing DP in these settings typically involves clipping client updates to a fixed sensitivity and adding Gaussian noise during aggregation, which helps protect participant-specific information while preserving overall learning utility. In 2021, the authors of [30] had observed a substantial increase in the use of DP in health research publications and, among others, the authors of [31, 32, 33, 34] developed a framework for privacy-preserving FL. In the context of the COVID-19 pandemic, the authors of [35] raised the awareness for a potential distortion of COVID-19 death rates when using DP with too small dataset groupings. During the COVID-19 pandemic, more than 1.9 million COVID-19 cases across Europe have been integrated in the DataSHIELD (Data Aggregation Through Anonymous Summary-statistics from Harmonised Individual-levEL Databases) platform [36], to allow for FL approaches to derive potential strategies against COVID-19 [37]. Among other things, DataSHIELD has already been used for other infectious diseases, e.g., to study malaria intervention in Mozambique [38].
In order to report dynamics and allow predictions by simultaneously addressing privacy concerns, COVID-19 case counts had been aggregated and reported on county-level throughout the pandemic in Germany [39]. We use these official aggregates as a proof-of-concept dataset to evaluate our method, which is intended for settings where finer-grained, potentially sensitive data remain with the LHAs. While aggregation reduces disclosure risk, it does not provide formal privacy guarantees [11]. Our work builds on established algorithms for federated learning with differential privacy [25, 28, 26] and selected DP-FL mechanism. To allow the reliable application in a public health context, we focus on a realistic infectious disease forecasting use case to study and evaluate client-level DP-FL on county-level COVID-19 data from two distinct epidemic phases in Germany. Selecting an established setup in which clients send clipped model updates and the server adds Gaussian noise calibrated to a target $(\varepsilon,\delta)$-DP guarantee via a Rényi DP accountant, we map out the privacy-utility tradeoff across a broad grid of privacy budgets and identify ranges of $\varepsilon$ for which forecasts remain practically useful. Our findings are intended to inform public health researchers and practitioners about which privacy budgets are compatible with accurate localized predictions.
@@
Our objective is time‐series forecasting of local case counts using data which is reported on local level. We construct supervised learning examples by using a historical window of size $H=10$ days over each administrative unit’s time series and predicting the case count $P=7$ days ahead. The window size $H$ and prediction horizon $P$ were determined empirically to ensure robust predictive performance. Let $c_{k,t}$ denote the raw, unnormalized number of daily reported cases for geographic unit $k$, $k=1,2,3,\ldots$ on day $t$. Therefore, the input vector
@@
is the case count $P$ days into the future. Missing reports are handled as zero reported cases. To mitigate weekly reporting artifacts (e.g., inconsistent weekend reporting), we applied a centered 7-day moving average. This smoothing is crucial not only for noise reduction but also to stabilize gradient updates in the FL process. For model evaluation, we construct examples by using a historical window of $H=10$ consecutive days and predicting the case count $P=7$ days ahead for each county. Within each epidemic period (November 2020 and March 2022) and for each county separately, we order all resulting input-label pairs by the date of the prediction target and assign the earliest $90~\%$ to the training set and the most recent $10~\%$ to the test set. In this way, all test samples correspond to later days than the training samples for the same county, reflecting the intended short-term forecasting use case. The November 2020 and March 2022 datasets are handled in separate experiments, i.e. models are trained and evaluated independently on each period. After splitting, we remove the location identifier from the feature vectors since they should not influence the model in any way.
@@
We use PyTorch [40] to build a fully connected multilayer perceptron (MLP) as our prediction model. The architecture is consistent across both spatial resolutions, community- and county-based; see the Results section for details on the datasets used. The input to the network is a vector of length $H$ containing the $H$ most recent daily case counts; with zero reported cases counting as daily case count. This vector is propagated through three fully connected hidden layers of sizes 128, 64, and 32 neurons, each followed by a rectified linear unit (ReLU) activation function. A final linear output unit produces the predicted case count $P$ days ahead. We use the mean squared error (MSE) as the loss function, which is minimized by the Adam optimizer with a learning rate of 0.001.
@@
We use a federated learning approach with client-level differential privacy. To formally define our privacy guarantees, we rely on the standard definition of $(\varepsilon,\delta)$-differential privacy [24]. In the following, we briefly summarize the background of differential privacy with our model choices.
Differential Privacy
Definition 1 (Differential Privacy [24])
A randomized mechanism $\mathcal{M}$} with range $\mathcal{R}$} satisfies $(\varepsilon,\delta)$-differential privacy, if for any two adjacent datasets $D$ and $D^{\prime}$, i.e., $D^{\prime}=D\cup\{x\}$ for some $x$ in the data domain (or vice versa), and for any subset of outputs $O\subseteq{\mathcal{R}}$, it holds that
$\Pr[{\mathcal{M}}(D)\in O]\leq e^{\varepsilon}\Pr[{\mathcal{M}}(D^{\prime})\in O]+\delta$ (3)
Intuitively, DP guarantees that an adversary, provided with the output of $\mathcal{M}$, can draw almost the same conclusions (up to $\varepsilon$ with probability larger than $1-\delta$) about any group no matter if it is included in the input of $\mathcal{M}$ or not [24]. This means, for any group owner, a privacy breach is unlikely to be due to its participation in the dataset.
We use the Gaussian mechanism to upper bound privacy leakage when transmitting information from clients to the server.
Definition 2
(Gaussian Mechanism [24]) Let $f:{\mathbb{R}}^{n}\rightarrow\mathbb{R}^{d}$ be an arbitrary function that maps $n$-dimensional input to $d$ logits with sensitivity being:
$S=\max_{D,D^{\prime}}\|f(D)-f(D^{\prime})\|_{2}$ (4)
over all adjacent datasets $D$ and $D^{\prime}\in{\mathcal{D}}$. The Gaussian Mechanism $\mathcal{M}_{\sigma}$, parameterized by $\sigma$, adds noise into the output, i.e.,
$\mathcal{M_{\sigma}}(x)=f(x)+\mathcal{N}(0,\sigma^{2}I).$ (5)
Adversarial Model. We assume an adversary that aims at inferring private information of the participating clients’ training datasets. We consider a strong adversary who can control a certain fraction of the participating clients. The adversary can access the training data, including all the parameters of the local models of the corrupted clients (i.e., it has white-box access to their local models). However, we consider that our server is trusted. The adversary is passive (i.e., honest-but-curious), that is, it follows the learning protocol faithfully.
Privacy Model. In federated learning, the notion of adjacent (neighboring) datasets used in DP generally refers to pairs of datasets differing by one client (client-level DP), or by one group of one user (group-level DP), or by one data point of one user (record-level DP). Our approach focuses on the client-level DP, where each client corresponds to a data holder, such as a Local Health Authority (LHA). Therefore, we aim to develop a solution that ensures client-level privacy. For instance, in scenarios involving collaborative LHAs, our goal is to safeguard any information unique to each LHA’s training data. The adversary should not be capable of discerning from the received model or its updates whether any client’s data has been included in the federated run (within $\varepsilon$ and $\delta$ bounds). We believe this adversarial model applies in many practical scenarios where private information extends across multiple samples within a client’s training data (for example, the presence of a group of samples, such as individuals with a specific feature). Differential Privacy provides plausible deniability not only for groups of a client’s samples but also for any client participating in the federated run. Consequently, any adverse privacy impact on clients (or their training samples) cannot be directly linked to their participation in the protocol run. The detailed privacy analysis, including the accounting via Rényi Differential Privacy (RDP), is provided in S1 Text..
DP-FL Algorithm
The core methodology employs FL with client-level DP, summarized in Algorithm 1. Our client-level DP-FL protocol is conceptually aligned with established DP-FL schemes [25, 26], where each selected client clips its local update to a fixed sensitivity, and the server adds Gaussian noise to the averaged update with a noise multiplier chosen via the Rényi DP accountant. Following standard practices in differentially private federated learning [31], the clipping threshold $S$ and the client sampling rate $q$ were determined empirically. All hyperparameters are additionally explained in Table 1. The training process basically follows four steps:
@@
Server:
Initialize common model ${\mathbf{w}}_{0}$
for $i=1$ to $T_{\mathsf{cl}}$ do
   Select a subset $\mathbb{K}$ of clients randomly with probability $q$
   Initialize list of updates $U=\mathbf{[}\,\mathbf{]}$
   for each client $k\in\mathbb{K}$ do
     $\Delta\widehat{{\mathbf{w}}}_{i}^{k}=\mathbf{Client}_{k}({\mathbf{w}}_{i-1})$
     Insert $\Delta\widehat{{\mathbf{w}}}_{i}^{k}$ into $U$
     end for
    $\Delta\overline{{\mathbf{w}}}_{i}=\frac{1}{m}\sum_{\Delta\widehat{{\mathbf{w}}}\in U}\Delta\widehat{{\mathbf{w}}}$
     if DP enabled then
       Calculate noise multiplier $c$ using an RDP accountant for $(\varepsilon,\delta,q,T_{\mathsf{cl}},E)$
       $\sigma=(S\cdot c)/m$
       Draw $\xi_{i}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$
       ${\mathbf{w}}_{i}={\mathbf{w}}_{i-1}+\Delta\overline{{\mathbf{w}}}_{i}+\xi_{i}$
       else
         ${\mathbf{w}}_{i}={\mathbf{w}}_{i-1}+\Delta\overline{{\mathbf{w}}}_{i}$
         end if
         end for
        $\mathbf{Client}_{k}({\mathbf{w_{i-1}}})$:
         $\widetilde{{\mathbf{w}}}_{i}={\mathbf{w}}_{i-1}$
         Train $\widetilde{{\mathbf{w}}}_{i}$ on local data $D_{k}$ for $E$ epochs using Adam ($\eta=0.001$)
         $\Delta{\mathbf{w}}_{i}^{k}=\widetilde{{\mathbf{w}}}_{i}-{\mathbf{w}}_{i-1}$
@@
         Output: $\Delta\widehat{{\mathbf{w}}}_{i}^{k}$
@@ Results
We use official reported county-level COVID-19 case data from the LHAs transferred to the Robert Koch Institute (RKI) [39] and accessed via the memilio-epidata Python package [42]. We consider two different pandemic stages, first, late-stage as of March 2022 and, second, early stage as of November 2020. Throughout this work, the November 2020 data and March 2022 data refer to two separate one-month windows of the county-level COVID-19 cases data. Models are trained and evaluated independently on each period. For each county and day, we obtain the newly confirmed infections plus estimated recoveries and deaths reported with the (assumed) day of infection. This data covers all 400 German counties. While the published data is stratified on county-level, the counties can, again, be divided by communities, summing up to a total of 10,786 communities in Germany. In Fig 1, Panel A, shows the population distribution across the counties, highlighting the variation in population sizes. The median county population is 147,524, while the mean population is 200,525, as several counties have a substantial larger population. The largest county by population has 3,685,265 inhabitants (Berlin, after merging all Berlin districts), and the total population across all counties is approximately 83.6 million. This heterogeneity in county sizes presents a particular challenge for our FL approach. Counties with larger populations typically report higher absolute case numbers. In addition, we show in Panel B the distribution of population sizes on community-level.
@@
Fig 2: Spatial resolution of German counties and communities and community-based case data. Panel A presents two maps shown side by side. The left map presents North Rhine-Westphalia (NRW) stratified into counties (thick black boundaries) and communities (thin gray boundaries), highlighting county ”Coesfeld” in red. The right map displays the stratification of Germany into federal states (thick black boundaries) and counties (thin gray boundaries), highlighting the federal state NRW in orange. Panel B shows the obtained community-based dataset for ”Coesfeld” for March 2022, with the county-level aggregate shown in red and individual community trajectories in various colors. The left map from Panel A is using geodata “Verwaltungsgebiete 1:250 000 (VG250)” from BKG (2026) dl-de/by-2-0, Data sources: https://sgx.geodatenzentrum.de/web_public/gdz/datenquellen/datenquellen_vg_nuts.pdf.
@@
To additionally analyze the impact of population heterogeneity, we stratified the test samples into five population clusters (0–50k, 50–100k, 100–200k, 200–500k, >>500k). Fig 6 shows the relative performance metrics (MAPE and MdAPE) for each cluster. While absolute errors naturally scale with population size, these relative metrics mainly remain stable across all clusters. This suggests that smaller counties are not substantially disadvantaged by the global model. Instead, they likely benefit from the shared learning of temporal patterns driven by data-rich counties, which stabilizes predictions on their potential sparser local data.
Fig 6: Prediction perfomance stratified by county population size. The panels show our metrics for different privacy levels ($\varepsilon\in\{0.3,0.5,1.0,2.0\}$ and non-DP) across five population clusters (0–50k, 50–100k, 100–200k, 200–500k, >>500k inhabitants) calculated from 15 runs, 75 federated rounds and 30 local epochs per round. The metrics include Mean Squared Error (MSE), Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and $R^{2}$ score. The shaded areas represent the standard deviation across all 15 runs.
@@
Privacy-utility tradeoff across epsilon values
To provide a more explicit view of the privacy-utility tradeoff, we conduct additional experiments on county-level data for both epidemic periods with a finer grid of privacy budgets,
$\varepsilon\in\{0.2,0.3,0.4,0.5,0.75,1.0,1.5,2.0,3.0,5.0,\infty\}$
We excluded $\varepsilon=0.1$, because the RDP accountant produces prohibitively large noise multipliers at this extreme privacy level, which leads to completely unstable training. For each value of $\varepsilon$ we repeated the DP-FL training $15$ times and recorded mean $\pm$ standard deviation of MSE, MAE, MAPE, and $R^{2}$. The resulting privacy-utility curves in Fig 8 plot these metrics as a function of the privacy budget on logarithmic axes for November 2020 and March 2022. For very strict privacy ($\varepsilon\leq 0.5$), the forecasts are essentially unusable, with MAPE values exceeding $1{,}000~\%$ and very large variability across runs. Between $\varepsilon=0.5$ and $\varepsilon=1.0$ the curves exhibit a sharp drop in all three error measures, reflecting the strongly nonlinear dependence of the RDP noise multiplier on $\varepsilon$. For moderately strong privacy ($\varepsilon\geq 2$), the DP-FL models in both periods closely approach the non-DP baseline and additional increases of $\varepsilon$ yield only marginal utility gains. The detailed summary statistics for all privacy budgets and both epidemic phases are reported in S1 Tab..
Fig 8: Privacy-utility tradeoff across privacy budgets. Privacy-utility curves for the November 2020 and March 2022 county-level datasets. Each panel shows mean $\pm$ standard deviation over $15$ runs of (A) MSE, (B) MAE, and (C) MAPE as a function of the privacy budget $\varepsilon$. Both axes are on a logarithmic scale, and the non-DP baseline is indicated at the right-hand side of each curve.
Convergence across federated rounds
To analyze how predictive performance evolves during federated training under different privacy guarantees, we track the test error as a function of the number of federated rounds. For each epidemic period (November 2020 and March 2022) and for privacy budgets $\varepsilon\in\{0.5,1.0,2.0\}$, as well as the non-DP baseline, we train the model for $75$ federated rounds with the same settings as in the main experiments (30 local epochs per round, clipping norm $S=0.5$, and $\delta=10^{-5}$). Every 5 rounds, we evaluate the current global model on the same test set and record the MAPE. We repeat this for 15 runs and report mean $\pm$ standard deviation across runs.
Fig 9 shows the resulting curves. For very strict privacy ($\varepsilon=0.5$), the test MAPE is already very high after a few rounds (around $80$-$100~\%$) and continues to increase over the course of training, with wide uncertainty bands that reflect large variability across runs. This indicates that the strong DP noise at this privacy level prevents stable learning and leads to increasingly noisy predictions. For intermediate privacy ($\varepsilon=1.0$), the MAPE initially stabilizes (November 2020) or decreases slightly (March 2022) in the early rounds, but then stays at a much higher level than for $\varepsilon=2.0$ (typically between $50$ and $150~\%$) and shows noticeable variation across runs. This suggests that additional rounds provide only small improvements and can even worsen performance when training is continued for too long.
In contrast, for $\varepsilon=2.0$ the curves quickly stabilize on a plateau with low MAPE values (about $25$-$30~\%$ in November 2020 and about $20$-$30~\%$ in March 2022). The non-DP baseline converges slightly faster and to slightly lower final MAPE values (around $25~\%$ in November 2020 and about $16$-$20~\%$ in March 2022), but the gap to $\varepsilon=2.0$ remains small in both epidemic phases. The small increase of the $\varepsilon=2.0$ curve at the last evaluation point in March 2022 lies well within the uncertainty variability band. Overall, these convergence patterns support the conclusion that privacy budgets around $\varepsilon\approx 2$ allow the model to reach useful forecasts, whereas very strict privacy ($\varepsilon=0.5$) leads to unstable training dynamics and unusable predictions.
Fig 9: Convergence of DP-FL across federated rounds for different privacy budgets. Test MAPE (%) as a function of the number of federated rounds for November 2020 (left) and March 2022 (right). Curves show mean $\pm$ standard deviation over 15 runs for privacy budgets $\varepsilon=0.5$ (blue), $\varepsilon=1.0$ (purple), $\varepsilon=2.0$ (orange), and Non-DP (red).